# AM-RADIO: Agglomerative Model – Reduce All Domains Into One If you can’t beat them, join them (and then beat them)

### v3

@@ 5.5 Semantic Segmentation Linear Probing
We train a linear head on top of the frozen features of the teachers and students alike and evaluate performance in the MMSeg [11] framework using the mIoU metric on ADE20k and PascalVOC2012 datasets. We use a training and evaluation crop size of 518x518 for models that support this resolution, or 224x224 for others. We use the “slide” evaluation mode with a stride of $\frac{2}{3}$ the crop size. We train the linear head for 160k steps using a total batch size of 16, a base learning rate of $10^{-3}$ and the AdamW optimizer.
@@ 5.6 Visual Question Answering
We replace the vision encoder in a LLaVA 1.5[35] setup with our own encoder. A 2-layer MLP is used to project frozen visual features into the language token space. Under the default LLaVA 1.5 settings, we pretrain a multimodal projection MLP and then run instruction tuning to finetune a Vicuna 7B-1.5 model[67]. We evaluate models using the validation sets of GQA [26], TextVQA [50] and ScienceQA [41] (image subset), and we score the model on the Test-Dev set of VQAv2 [20] using EvalAI[60]. We use an input image resolution of 336x336 when the vision encoder model supports it, and fall back to 224x224 otherwise. Experimental results are compiled in Table 7. Qualitative results are shown in Figure 3.
6 Key Insights
Outperforming teachers. Most Vision Foundation Models (VFMs) have unique properties such as language grounding (CLIP), dense correspondences (DINOv2), and detailed segmentation (SAM). Distillation allows uniting all these properties in a single model that even outperforms each of them. We have also observed that better teachers yield better students, which allows RADIO to absorb the current SOTA foundation models at a given point in time.
Training efficiency. Our distillation framework consumes only 2%-5% of the data samples compared to training CLIP [18] or DINO from scratch. Even with the costs of running multiple teachers during training, the overall compute is only a quarter of that required for training CLIP from scratch.
Feature distillation loss. We observe the crucial importance of full feature distillation to boost the performance of the teacher in dense image understanding tasks, such as an 18% relative improvement on ADE20K. Future research focusing on developing new loss formulations could further enhance performance.
@@ 6 Key Insights
Efficient backbone. Most of the existing VFMs use ViTs as the model architecture for the vision encoder. Based on our analysis of distilling efficient backbones, we conclude that most model designs are overly tailored towards supervised training on ImageNet1K and, as a result, do not scale well to VFM settings due to the low feature resolution at the output. We designed a new vision backbone, E-RADIO, with a hybrid CNN-Transformer architecture that significantly outperforms the competition. While it is at least 6x faster, it still lags behind pure transformers on dense tasks.
Conclusions
We propose a multi-teacher distillation framework to efficiently train vision foundation models (VFMs). We demonstrate that it can unify unique properties of other VFMs such as CLIP, DINOv2, and SAM. Our models outperform the original teachers in multiple tasks under the same model architectures while exhibiting greater capability. On the efficiency side, we designed an architecture, E-RADIO, that is up to 10x faster, while closely matching or even outperforming the teachers. We believe our work will advance the field of efficient VFM design and benchmarking.
Acknowledgment
We would like to thank Hongxu (Danny) Yin, Karan Sapra, and Francesco Ferroni for their important feedback on the project and Philipp Fischer and Lukas Vögtle for help with training infrastructure.
@@ Acknowledgment
Appendix A Teacher Comparison
Family Arch Throughput Zero-Shot k-NN ADE20k VOC GQA SQA TextVQA VQAv2
OpenCLIP [9] ViT-H/14 556 77.19 81.10 40.04 68.03 57.86 67.53 48.11
MetaCLIP [59] ViT-H/14 556 80.51 82.12 35.39 62.62 67.85 68.77 52.02 75.46
DINOv2 [43] ViT-L/14 1126 N/A 83.44 46.55 82.07 70.16 63.76 45.49 75.37
ViT-G/14 344 N/A 83.41 47.53 82.95 69.29 64.60 45.25 74.56
SAM [31] ViTDet-L/16 26 N/A 20.41 26.56 34.34 59.87 67.08 43.90 62.46
ViTDet-H/16 16 N/A 22.39 28.17 32.16 57.62 63.26 43.34 60.00
Table 10: Comparison of teachers models. ”Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K and VOC refer to linear probe semantic segmentation mIoU on the ADE20K [68] and PascalVOC2012 validation datasets. Note: for SAM, throughput and all other metrics were measured using the only supported image resolution of 1024x1024. We report k-NN metrics and Visual Question Answering average pooling the 64x64 patch tokens into a 16x16 feature map, and emphasize that SAM was not designed for these tasks. For other models, we used an image resolution of 224x224 for image classification, 336x336 for visual QA and 518x518 for semantic segmentation.
Figure 4: Comparison of teachers within our suite of downstream evaluation tasks.
Teacher metrics are summarized in Table 10. Interestingly, we note that even though MetaCLIP performs better than OpenCLIP on image-level classification and visual question answering metrics, it exhibits lower performance on semantic segmentation.
Appendix B E-RADIO architecture details
@@
B.1 Multi-Resolution Attention
@@ Appendix C ViTDet Augmentation
Appendix D Automatic Loss Balancing
D.1 Uncertainty
@@
The function $\psi(x)$ is the familiar ”softplus” nonlinear activation function. We drop the division by 2 on the left because, assuming outputs are initially $b\sim\mathcal{N}(0,\sigma^{2})$, then the loss weights will initially have an expected value of 1, matching the naive weighting. On the right, we replace $\frac{b_{k}}{2}$ with $\psi(x)$ for a few reasons:
@@ D.2 AdaLoss
Appendix E Visual Question Answering Samples
Figure 7: Visualization of the LLaVA attention maps over the visual features produced by a RADIO encoder. We use one sample image from the GQA[26] validation set and one associated question: ”What color is the helmet in the middle of the image?”. For each layer in the language model, we retrieve attention scores for all positions of the visual tokens, average them over all attention heads, and overlay corresponding heat maps with the input image. We can see that as we progress through the layers, the model’s attention focuses on the relevant part of the image. The model’s answer is ”Blue”.
@@ Appendix E Visual Question Answering Samples
Figures 8 to 13 show sample questions from our Visual Question Answering datasets, together with sample answers when using our vision encoders in a LLaVA setup.
Appendix F Original Model Comparison Table
Table 7 shows updated RADIO results after the original publication date. For reference, we provide the original table as table 11.
Family Arch Throughput Zero-Shot k-NN ADE20k VOC GQA SQA TextVQA VQAv2
OpenCLIP [9] ViT-H/14 556 77.19 81.10 40.04 68.03 57.86 67.53 48.11 67.31
MetaCLIP [59] ViT-H/14 556 80.51 82.12 35.39 62.62 67.85 68.77 52.02 75.46
OpenAI CLIP [46] ViT-L/14 1126 75.54 79.80 36.51 67.04 68.80 67.03 53.25 76.47
DINOv2 [43] ViT-L/14 1126 N/A 83.44 46.55 82.07 70.16 63.76 45.49 75.37
ViT-g/14 344 N/A 83.41 47.53 82.95 69.29 64.60 45.25 74.56
SAM [31] ViTDet-L/16 14 N/A 20.41 26.56 34.34 61.52 63.36 43.29 64.91
RADIO (Ours) E-RADIO-L 3697 77.87 83.73 45.50 79.95 66.22 65.54 43.51 70.74
ViT-H/14 556 78.62 84.17 49.01 83.51 70.46 65.89 53.36 77.66
Table 11: Comparison of teachers and RADIO models. ”Zero-Shot” and k-NN are computed on ImageNet-1K. ADE20K and VOC refer to linear probe semantic segmentation mIOU on the ADE20K [68] and PascalVOC2012 validation datasets. RADIO outperforms all teacher models on most metrics, and then E-RADIO-L is about 6.6x faster than ViT-H/14 while attaining similar metrics on certain tasks.

### v4

@@ 5.3 Semantic Segmentation Linear Probing
We train a linear head on top of the frozen features of the teachers and students alike and evaluate performance in the MMSeg [12] framework using the mIoU metric on ADE20k and PascalVOC2012 datasets. We use a training and evaluation crop size of 512 for RADIO, 518 for DINOv2, and the native resolution for the others. We use the “slide” evaluation mode with a stride of $\frac{2}{3}$ the crop size. We train the linear head for 160k steps using a total batch size of 16, a base learning rate of $10^{-3}$ and the AdamW optimizer.
@@ 5.4 Visual Question Answering
We replace the vision encoder in a LLaVA 1.5[37] setup with our own encoder. A 2-layer MLP is used to project frozen visual features into the language token space. Under the default LLaVA 1.5 settings, we pretrain a multimodal projection MLP and then run instruction tuning to finetune a Vicuna 7B-1.5 model[71]. We evaluate models using the validation sets of GQA [27], TextVQA [52], POPE [36] (popular), and we score the model on the Test-Dev set of VQAv2 [21] using EvalAI[62]. We use the vision encoder’s native input resolution, resizing the long edge and padding the short edge. Experimental results are compiled in Table 1. Owing to the increased input resolution flexibility of RADIO, we resize the long edge of the image to 432px aspect preserving, only padding to the nearest multiple of the patch size. This results in 462462~{}462462 tokens on average, versus the $576$ tokens required by the 336px patch-14 encoders, a 20% reduction.
6 Conclusion and Key Insights
Figure 4: RADIO “mode switches” when resolution is increased. In the plot, we show the MSE error between the RADIO features coming from its DINOv2 head at different resolutions, versus the features actually produced by DINOv2 at 518px. We bilinearly interpolate the RADIO features to match the DINOv2 feature resolution. At 720px, there is a sudden jump in the error, which corresponds with a complete change in color space in the image.
Most VFMs have unique properties such as language grounding (CLIP), dense correspondences (DINOv2), and detailed segmentation (SAM), but also large holes in capability. Distillation allows uniting all these properties in a single model that often outperforms any of the teachers. We have also observed that better teachers yield better students, which allows RADIO to absorb and challenge the current SOTA foundation models at a given point in time.
Feature distillation loss. We observe the crucial importance of full feature distillation to boost the performance of the teacher in dense image understanding tasks, such as an 18% relative improvement on ADE20K.
@@ 6 Conclusion and Key Insights
Dense features. As seen in figure 1, RADIO is capable of producing high resolution and low-noise features. Given the adaptor heads, it is able to reasonably reproduce the features generated by the teachers, and the backbone features themselves are highly coherent. An issue we identified, however, shown in figure 4 is that RADIO appears to have a latent ‘low resolution’ and ‘high resolution’ mode, likely due to the partitioned training between CLIP+DINO and SAM objectives, which we intend to fix in future work.
Efficient backbone. Based on our analysis of distilling efficient backbones, we conclude that most model designs are overly tailored towards supervised training on ImageNet1K, and as a result, do not scale well to VFM settings. We designed a new vision backbone, E-RADIO, with a hybrid CNN-Transformer architecture that improves upon the Pareto frontier.
@@
Appendix A E-RADIO architecture details
@@
A.1 Multi-Resolution Attention
@@ A.2 Configurations
Appendix B PCA Visualizations
We visualize various models using PCA to reduce the model’s spatial feature dimensionality down to 3 dimensions, and directly map those to RGB. Most models are only able to handle square inputs at fixed resolutions, however DINOv2 and RADIO can handle arbitrary resolutions and aspect ratios, so we visualize them in both settings.
B.1 Square Models
Model Resolution Images
OpenCLIP-H/14 224
MetaCLIP-H/14 224
SigLIP-M/14 384
InternViT-6B 224
448
DFN CLIP 378
OpenAI CLIP 336
DINOv2-g 518
SAM-H 1024
RADIO 512
1024
B.2 Flexible Models
Model Resolution Images
DINOv2-g 518
1022
2044
RADIO 512
1024
2048
@@ Appendix C ViTDet Augmentation
Appendix D Comparison with SAM-CLIP [57]
Concurrently with our work, SAM-CLIP was introduced as a method of fusing SAM and CLIP into a single model. Due to the concurrency of effort, we don’t compare our model with the full suite of metrics demonstrated in their method, however, we do have some overlap in key metrics such as Zero-Shot ImageNet-1k, and ADE20k semantic segmentation via linear probing. We present the comparison in table 9, however we note that there are enough differences between these two models that we can’t conclude one way or another what is the superior approach. Instead we’ll argue that DINOv2 does a better job of ADE20k linear probing than SAM, and thus our significantly higher quality on this metric is likely due to the inclusion of DINOv2, which is a key introduction with our approach.
Family Model Zero-Shot ADE20k
SAM ViTDet-H/16 28.2
DFN CLIP ViT-H/14 83.9 31.7
SAM-CLIP ViTDet-B/16 71.7 38.4
RADIO ViT-H/14 82.7 51.3
Table 9: We compare our common key metrics with those demonstrated in SAM-CLIP [57]. We note that there are numerous differences between the two approaches, including model capacity and architecture. SAM-CLIP uses the ViT-B variant of SAM as a starting point, which implies it’s a ViTDet-B/16 architecture. As a result of this choice, their metrics are computed at a resolution of 1024. RADIO trains a vanilla ViT-H/14 from scratch, and as a result of the flexibility gained via the CPE method, we evaluate Zero-Shot ImageNet1k at a resolution of 432, and we run ADE20k linear probing at a resolution of 512 using the exact same weights. We note that Zero-Shot quality is largely determined by the quality of the CLIP teacher and the capacity of the student. We attribute our superior quality on ADE20k semantic segmentation largely to our inclusion of DINOv2 as a teacher.
Appendix E Automatic Loss Balancing
E.1 Uncertainty
@@
The function $\psi(x)$ is the familiar “softplus” nonlinear activation function. We drop the division by 2 on the left because, assuming outputs are initially $b\sim\mathcal{N}(0,\sigma^{2})$, then the loss weights will initially have an expected value of 1, matching the naive weighting. On the right, we replace $\frac{b_{k}}{2}$ with $\psi(x)$ for a few reasons:
@@ E.2 AdaLoss
Appendix F Visual Question Answering Samples
Figure 7: Visualization of the LLaVA attention maps over the visual features produced by a RADIO encoder. We use one sample image from the GQA[27] validation set and one associated question: "What color is the helmet in the middle of the image?". For each layer in the language model, we retrieve attention scores for all positions of the visual tokens, average them over all attention heads, and overlay corresponding heat maps with the input image. We can see that as we progress through the layers, the model’s attention focuses on the relevant part of the image. The model’s answer is "Blue".
@@ Appendix F Visual Question Answering Samples
Figures 8 to 12 show sample questions from our Visual Question Answering datasets, together with sample answers when using our vision encoders in a LLaVA setup.